OpenStudy (anonymous):

Solve the equation x +1 − 2 = x − 7 for x and check your answer(s)

OpenStudy (mathstudent55):

First combine like terms on both sides if possible. Like terms are terms that have exactly the same variables. Terms with x are like terms and can be combined together, and terms without x are like terms and can be combined together. In this case you have two numbers on the left side that can be combined togerther x + 1 - 2 = x - 7 Combine the numbers on left side: x - 1 = x - 7 Now to get all x's on left side and all numbers on right side, subtract x from both sides and add 1 to both sides x - 1 = x - 7 -x + 1 = -x + 7 ------------------ 0 + 0 = 0 + 7 You get 0 = 7. Since this is a false statement, this equation has no solution.

OpenStudy (anonymous):

oo my you know what lol im so srry I wrote this wrong cause I was looking at what I did in my notes and what u were writing and im like?? it should be \[\sqrt{x+1}-2=x-7\]

OpenStudy (mathstudent55):

To solve this type of of equation is isolate the square root on one side, then square both sides.

OpenStudy (mathstudent55):